A glass rod transferred 6.242x10^11 electrons to an electroscope. What is the charge on the electroscope?
Answer #1

Charge of 1 electron = 1.6*10^-19 C

Charge on 6.242x10^11electrons is 6.242x10^11 *1.6*10^-19 C = 9.9*10^-8 C

charge on the electroscope is 9.9*10^-8 C